OpenStudy (anonymous):

2x+7=x+3 SHOW ALL STEPS

OpenStudy (anonymous):

So 2x+7=x+3 Mivr x to other side x+7=3 Subtract7 X=-4

rebeccaxhawaii (rebeccaxhawaii):

add like terms

OpenStudy (serenity74):

2x + 7 = x+3 2x - x = 3-7 x = -4

rebeccaxhawaii (rebeccaxhawaii):

@you.mad.bro please do not give the answer let the person with the question figure it out

OpenStudy (serenity74):

You bring the x-terms to one side and the variables to one side and then solve for x.

OpenStudy (anonymous):

Ik this sweetie

OpenStudy (serenity74):

^ chill kid back off

OpenStudy (anonymous):

Lol I'm 18

OpenStudy (serenity74):

Does it look like I care? Clearly not. Follow the COC. Age doesn't matter. So chill.

OpenStudy (anonymous):

Yes ik I read the coc silly I want trying to give direct

OpenStudy (serenity74):

*Wasn't Please close this question, thank you! Have a lovely day! :)